OpenStudy (anonymous):

find the equation of the line useing the following info show all your work (-6,-3)(5,8)

OpenStudy (anonymous):

use the formula:

OpenStudy (anonymous):

(y-ysub1)=(ysub2-ysub1/xsub2-xsub1)*(x-xsub1) (y-(-3))=(8-(-3)/5-(-6)) (x-(-6)) y+3=(11/11)(x+6) y+3=(x+6) x-y=-3

OpenStudy (anonymous):

you can check it by substituting the values or coordinates